# Similarity-Dissimilarity Loss for Multi-label Supervised Contrastive Learning

### v4

@@ 2.5 Case Analysis
[…]
Applying these values to Eq. (14) and (15), we obtain:
[…]
@@
When evaluating Eq. (13), these distinct relations ($R2$ through $R5$) generate unique loss values, establishing the following inequalities:
@@
The proposed loss function effectively discriminates among the five distinct relations through a principled re-weighting mechanism, as formulated in Eq. (13), (14), and (15), comparing to existing methods in MSCL.
@@ 2.6 Theoretical Analysis
The proposed loss function incorporates a weighting mechanism through the product of factors $\mathcal{K}_{i,p}^{s}$ and $\mathcal{K}_{i,p}^{d}$. By construction, the similarity-dissimilarity factor $\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}$ is constrained to the closed interval $[0,1]$ across all possible relational configurations. Hence, it is written as:
@@
See Appendix A.3. ∎
Based on Theorem 1, 2, and 3, the product of weighting factors $\mathcal{K}_{i,p}^{s}$ and $\mathcal{K}_{i,p}^{d}$ is bounded within the interval $[0,1]$, which aligns with fundamental principles of loss functions and set-theoretic relations. The non-negative lower bound adheres to the essential property of loss functions being strictly positive [8]. Given that our proposed loss function generalizes the supervised contrastive loss [1] and incorporates multi-label relation definitions, the upper bound naturally equals $1$. Furthermore, this mathematical framework demonstrates that our proposed contrastive loss can dynamically adjust the weighting factor within $[0,1]$, effectively differentiating sample features with rigorous mathematical justification for both the formulation and efficacy of the loss function.
@@
$|\tilde{\boldsymbol{y}}_{p_{3}}|=|\tilde{\boldsymbol{y}}_{p_{4}}|$ (22)
@@
$|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|\leq|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|$ (24)
@@
See Appendix A.5. ∎
Theorem 4 and 5 establish strict dominance relations between relation types $R3$, $R4$, and $R5$, demonstrating that $\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ when $|\tilde{\boldsymbol{y}}_{p_{3}}|=|\tilde{\boldsymbol{y}}_{p_{4}}|$ and $\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ when $|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|\leq|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|$. These inequalities, proved through careful mathematical derivation using set cardinality properties and fundamental principles of real analysis, reveal a well-defined hierarchical structure in the weighting factors. This hierarchical relations ensures that our loss function appropriately modulates the contribution of different relation types during the learning process, providing theoretical guarantees for the effectiveness of our proposed approach in capturing complex relations within the data.
@@
3 Experiments
We conduct the experiments to compare Similarity-Dissimilarity Loss with current contrastive loss functions (ALL, ANY, and MulSupCon) in a comprehensive evaluation, considering: (i) Data modality: image and text data; (ii) Domain-specific: general text data (AAPD) and medical domain (MIMIC III and IV); (iii) Data distribution: full setting (extreme long-tailed distribution) and top-50 frequent labels setting; (iv) ICD code versions: ICD-9 and ICD-10, and (v) Models: ResNet-50, RoBERTa-based, Llama-3.1-8B, and PLM-ICD. For comprehensive information, please refer to Appendix B for detailed dataset descriptions and evaluation metrics, Appendix C for baseline models and encoder architectures, and Appendix D for implementation details.
4 Results and Analysis
More experiment results for the evaluation on medical domain in Appendix E
@@ 4.1 Evaluation on Image
Figure 2(a) illustrates the comparison between Similarity-Dissimilarity Loss and MulSupCon as measured by micro- and macro-F1 metrics. The results indicate that our method yields substantially greater improvements in macro-F1 compared to micro-F1 across all image datasets. Specifically, macro-F1 increases by $5.7\%$ on MS-COCO and $5.8\%$ on NUS-WIDE, whereas micro-F1 exhibits more modest improvements of $2.9\%$ and $2.0\%$, respectively. Macro-F1 assigns equal importance to each class regardless of its frequency, rendering it particularly appropriate for evaluating performance on imbalanced datasets where minority class prediction accuracy is critical [8, 7]. In contrast, micro-F1 places more considerable weight on classes with more samples, making it more appropriate when larger classes should have a more potent influence on the overall score [8, 9]. Multi-label classification inherently faces more pronounced challenges with long-tailed distributions than single-label classification due to exponential output space complexity, intricate label co-occurrence patterns, and high annotation costs [7]. The observed superior improvement in macro-F1 metrics provides compelling evidence that our method demonstrates exceptional efficacy in addressing long-tailed distribution challenges, a capability particularly crucial in multi-label scenarios.
However, on the PASCAL dataset, our method demonstrates mere marginal improvements, with gains of $0.88/0.84/0.17$ in micro/macro-F1/mAP, respectively. This limited enhancement can be attributed to the structural characteristics of PASCAL, wherein the average number of labels per instance is approximately $1.5$ (as detailed in Table 3), causing the task to approximate single-label classification, particularly when the batch size is limited [1]. Consequently, loss functions specifically designed for multi-label scenarios exert minimal influence on model performance under these conditions. As Audibert et al. [10] have demonstrated, the cardinality of the label space constitutes a significant determinant of model efficacy within MSCL .
Furthermore, Figure 2(b) reveals that the standard deviation across four methods for PASCAL equals $0.58/0.64/0.27$ in micro/macro-F1/mAP, which are considerably lower than the corresponding standard deviations observed for the MS-COCO and NUS-WIDE. This statistical finding suggests that the efficacy of specialized multi-label loss functions diminishes significantly when the average label cardinality per instance approaches $1$ in MSCL. This finding further corroborates our theoretical analysis and hypothesis in the Section 2, wherein Similarity-Dissimilarity Loss degenerates to single-label scenarios (see Eq. (17)).
@@
4.2 Evaluation on Text
We further evaluate our method on general text data, and the results demonstrate that our proposed loss function consistently surpasses baseline methods for both RoBERTa and Llama models across all metrics on the AAPD dataset (See Table 2). In contrast to the significant performance gains observed on image data, Similarity-Dissimilarity Loss achieves more modest enhancements of $0.90/1.79$ in micro/macro-F1 scores on RoBERTa, and $0.89/1.84$ on Llama. This attenuated performance differential can be attributed to the extensive knowledge already encoded within LLMs through their comprehensive pre-training paradigms [11].
Method RoBERTa Llama
Micro-F1 Macro-F1 Micro-F1 Macro-F1
ALL 73.23 59.41 74.32 60.47
ANY 72.31 58.55 73.41 59.63
MulSupCon 73.64 60.52 74.72 61.58
Ours 74.54 62.31 75.61 63.42
@@ 5 Conclusion
In this paper, we introduce multi-label relations and propose a Similarity-Dissimilarity Loss for MSCL, which dynamically re-weights the loss by the similarity-dissimilarity factor. We provide theoretical grounded proof for our method through rigorous mathematical analysis that supports the formulation and effectiveness of the proposed loss function. The experimental results further confirm the effectiveness and robustness of our method in image, text and medical domain. Despite the advancements, limitations warrant acknowledgment. The Similarity-Dissimilarity Loss, while effective via linear factors in MSCL, may benefit from further explore non-linear similarity-dissimilarity weighting factors and learnable hyperparameter for contrastive loss function to capture more complex semantic relations and long-tailed distribution.
Appendix A Theorem Proofs
A.1 Proof of Theorem 1
@@
$|\tilde{\boldsymbol{y}}_{p}^{s}|$ $=|\mathcal{S}\cap\mathcal{T}|=|\varnothing|=0$
$\therefore\mathcal{K}_{1}^{s}$ $=\frac{|\tilde{\boldsymbol{y}}_{p}^{s}|}{|\tilde{\boldsymbol{y}}_{i}|}=\frac{0}{|\tilde{\boldsymbol{y}}_{i}|}=0$
@@ A.3 Proof of Theorem 3
We first establish the strict positivity. Given $|\mathcal{S}|>0$ and conditions (28)-(29), we have:
@@
Case 1 ($m=3$): By Eq. (10), we have three conditions: $\mathcal{S}\cap\mathcal{T}\neq\varnothing$, $\mathcal{S}\nsubseteq\mathcal{T}$, and $\mathcal{T}\nsubseteq\mathcal{S}$. These conditions lead to:
@@
Case 2 ($m=4$): When $m=4$, by Eq. (11), we have $\mathcal{S}\supseteq\mathcal{T}$. This subset relation implies:
@@
Case 3 ($m=5$): When $m=5$, by Eq. (12), we have $\mathcal{S}\subsetneq\mathcal{T}$. This subset relation implies:

### v5

@@ 2.5 Case Analysis
$\left\{\begin{aligned} |\tilde{\bm{y}}_{p_{j}}^{s}|&=\{0,3,1,2,3\}_{j=1,2,3,4,5}\\ |\tilde{\bm{y}}_{p_{j}}^{d}|&=\{3,0,2,0,2\}_{j=1,2,3,4,5}\end{aligned}\right$.
Applying these values to Equation 14 and 15, we obtain:
$\left\{\begin{aligned} \mathcal{K}_{i,p}^{s}&=\{0,1,\frac{1}{3},\frac{2}{3},1\}\\ \mathcal{K}_{i,p}^{d}&=\{\frac{1}{4},1,\frac{1}{3},1,\frac{1}{3}\}\end{aligned}\right$.
@@
When evaluating Equation 13, these distinct relations ($R2$ through $R5$) generate unique loss values, establishing the following inequalities:
@@
The proposed loss function effectively discriminates among the five distinct relations through a principled re-weighting mechanism, as formulated in Equation 13, 14, and 15, comparing to existing methods in MSCL.
@@ 2.6 Theoretical Analysis
The proposed loss function incorporates a weighting mechanism through the product of factors $\mathcal{K}_{i,p}^{s}$ and $\mathcal{K}_{i,p}^{d}$. By construction, the similarity-dissimilarity factor $\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}$ is constrained to the closed interval $[0,1]$ (i.e., the upper and lower bounds) across all possible relational configurations. Hence, it is written as:
@@
See Appendix B.3. ∎
Based on Theorem 1, 2, and 3, the product of weighting factors $\mathcal{K}_{i,p}^{s}$ and $\mathcal{K}_{i,p}^{d}$ is bounded within the interval $[0,1]$, which aligns with fundamental principles of loss functions and set-theoretic relations. The non-negative lower bound adheres to the essential property of loss functions being strictly positive (LeCun et al., 2015). Given that our proposed loss function generalizes the supervised contrastive loss (Khosla et al., 2020) and incorporates multi-label relation definitions, the upper bound naturally equals $1$. Furthermore, this mathematical framework demonstrates that our proposed contrastive loss can dynamically adjust the weighting factor within $[0,1]$, effectively differentiating sample features with rigorous mathematical justification for both the formulation and efficacy of the loss function.
@@
$|\tilde{\bm{y}}_{p_{3}}|=|\tilde{\bm{y}}_{p_{4}}|$ (22)
@@
$|\tilde{\bm{y}}_{p_{5}}^{d}|\leq|\tilde{\bm{y}}_{p_{3}}^{d}|$ (24)
@@
See Appendix B.5. ∎
Theorem 4 and 5 establish strict dominance relations between relation types $R3$, $R4$, and $R5$, demonstrating that $\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ when $|\tilde{\bm{y}}_{p_{3}}|=|\tilde{\bm{y}}_{p_{4}}|$ and $\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ when $|\tilde{\bm{y}}_{p_{5}}^{d}|\leq|\tilde{\bm{y}}_{p_{3}}^{d}|$. These inequalities, proved through careful mathematical derivation using set cardinality properties and fundamental principles of real analysis, reveal a well-defined hierarchical structure in the weighting factors. This hierarchical relations ensures that our loss function appropriately modulates the contribution of different relation types during the learning process, providing theoretical guarantees for the effectiveness of our proposed approach in capturing complex relations within the data.
@@
2.7 Re-weighting Mechanism
The proposed re-weighting mechanism is designed to assign soft weights to positive pairs in multi-label contrastive learning, ensuring that the contribution of each positive to the objective function reflects the degree of semantic alignment with the anchor. Formally, the weighting factor for a positive sample $p$ associated with anchor $i$ is given by the product
$w_{i,p}\;=\;\mathcal{K}_{i,p}^{s}\,\mathcal{K}_{i,p}^{d},\quad w_{i,p}\in[0,1],$
where $\mathcal{K}_{i,p}^{s}$ captures similarity based on shared labels, and $\mathcal{K}_{i,p}^{d}$ penalizes mismatches due to additional labels.
The similarity component $\mathcal{K}_{i,p}^{s}$ is monotonically increasing with respect to the number of shared labels between $i$ and $p$. Let $L(i)$ denote the set of labels associated with sample $i$. Then,
$\mathcal{K}_{i,p}^{s}\;=\;f\big(|L(i)\cap L(p)|\big),$
where $f(\cdot)$ is a non-decreasing mapping that promotes stronger alignment when $p$ and $i$ share more labels, thereby reflecting higher semantic agreement.
The dissimilarity component $\mathcal{K}_{i,p}^{d}$ instead captures the penalty incurred by extraneous labels present in $p$ but absent in $i$. Specifically,
$\mathcal{K}_{i,p}^{d}\;=\;g\big(|L(p)\setminus L(i)|\big),$
where $g(\cdot)$ is a monotonically decreasing function, ensuring that positives with many additional labels contribute less, since these labels may correspond to irrelevant or noisy semantics.
The resulting product $w_{i,p}$ quantifies the relational quality between $i$ and $p$. Several important cases arise:
  1. 1.
    Perfect alignment ($R2$): If $L(i)=L(p)$, then $\mathcal{K}_{i,p}^{s}$ is maximized and $\mathcal{K}_{i,p}^{d}=1$, yielding $w_{i,p}\approx 1$.
  2. 2.
    Partial overlap (R3R3R5R5): If $L(i)\cap L(p)\neq\varnothing$ but $L(i)\neq L(p)$, then $\mathcal{K}_{i,p}^{s}$ is non-zero but $\mathcal{K}_{i,p}^{d}<1$, leading to an intermediate weight $0<w_{i,p}<1$.
  3. 3.
    Disjoint labels ($R1$): If $L(i)\cap L(p)=\varnothing$, then $\mathcal{K}_{i,p}^{s}=0$ and $w_{i,p}=0$, ensuring no contribution to the loss.
This formulation provides a principled mechanism for modulating the contrastive signal. High-quality positives, characterized by substantial semantic overlap and minimal extraneous labels, are emphasized in the embedding space. Conversely, positives with excessive label mismatches are downweighted, thereby mitigating the risk of noisy or misleading training signals. Importantly, the continuous nature of $w_{i,p}$ allows the model to smoothly calibrate the pulling force across a spectrum of semantic relationships, rather than making binary inclusion–exclusion decisions.
Such a mechanism is particularly critical under long-tailed and ambiguous multi-label distributions, where the frequency of classes and the degree of label co-occurrence vary significantly. By balancing similarity and mismatch in a mathematically grounded manner, the re-weighting scheme ensures that the learned representations remain both semantically faithful and robust to noise.
3 Experiments and Results
We conduct the experiments to compare Similarity-Dissimilarity Loss with current contrastive loss functions (ALL, ANY, and MulSupCon) in a comprehensive evaluation, considering: (i) Data modality: image and text data; (ii) Domain-specific: general text data (AAPD) and medical domain (MIMIC III and IV); (iii) Data distribution: full setting (extreme long-tailed distribution) and top-50 frequent labels setting; (iv) ICD code versions: ICD-9 and ICD-10, and (v) Models: ResNet-50, RoBERTa-based, Llama-3.1-8B, and PLM-ICD. For comprehensive information, please refer to computing cost analysis in Appendix C, detailed dataset descriptions and evaluation metrics in Appendix D, baseline models and encoder architectures in Appendix E, and implementation details in Appendix F
@@ 3.1 Evaluation on Image
The experimental results in Table 1 demonstrate that our proposed loss function outperforms baselines across all metrics on image datasets. The substantial improvements in macro-F1 (Figure 2(a)) provide compelling evidence that our method demonstrates exceptional efficacy in addressing long-tailed distribution challenges, a capability particularly crucial in multi-label scenarios. However, on the PASCAL dataset, the method yields only marginal gains due to its low average label cardinality (approximately 1.5), which causes the task to approximate single-label classification; consequently, multi-label loss functions exert limited influence, consistent with prior findings that label cardinality is a significant determinant of MSCL performance (Audibert et al., 2024). Moreover, Figure 2(b) shows that PASCAL exhibits much lower variance across methods compared to MS-COCO and NUS-WIDE, indicating that specialized multi-label loss functions become less effective when label cardinality approaches one. This observation supports the theoretical analysis in Section 2, where the Similarity–Dissimilarity Loss reduces to single-label cases (Equation 17). A detailed analysis and discussion the results on image data is provided in Appendix G.1
@@
3.2 Evaluation on Text and Medical Domain
We conduct comprehensive evaluations on several text datasets, including AAPD, MIMIC-III, and MIMIC-IV, with the MIMIC datasets belonging to the medical domain. Detailed results and analyses on AAPD are provided in Appendix G.2, while the extensive evaluation and discussion of the medical domain are presented in Appendix G.3. Furthermore, our proposed method achieves state-of-the-art performance on MIMIC-III-Full by employing the MSCL framework with Similarity–Dissimilarity Loss and the PLM-ICD encoder (see Table 6 and Appendix G.3).
@@ 4 Conclusion
In this paper, by systematically formulating multi-label relations, we establish a principled framework for sample identification in MSCL. Building on this foundation, we propose the Similarity–Dissimilarity Loss, which dynamically re-weights samples based on similarity and dissimilarity factors, supported by rigorous theoretical analysis that establishes its mathematical soundness and generalizability to both single-label and multi-label settings. Extensive experiments across image, text, and medical domains demonstrate that our method consistently outperforms strong baselines, achieving state-of-the-art results on the MIMIC datasets. These findings highlight the effectiveness, robustness, and broad applicability of the proposed approach, offering a unified paradigm that advances the study and practice of contrastive learning in complex multi-label scenarios. Moreover, the limitations of this work is discussed in Appendix I.
Ethics statement
We use the publicly available clinical dataset MIMIC-III and IV, which contain de-identifed patient information. We do not see any ethics issues here in this paper.
Reproducibility statement
Implementation details for reproducing the experimental results are provided in Appendix F, while the data description and access information are presented in Appendix D.
Appendix
Appendix A Further Explanation of Similarity-Dissimilarity Loss
A.1 Explanation of Scaling Temperature
We now explain why we introduce the similarity–dissimilarity weighting by multiplying the factor $\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}$ directly into the numerator of the log-softmax loss (Equation 13), rather than scaling the temperature parameter $\tau$ as the formulation is given by
$\mathcal{L}_{i}^{\text{our}}=\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\exp\!\left((\bm{z}_{i}\cdot\bm{z}_{p})/\tau\cdot\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}\right)}{\sum_{a\in\mathcal{A}(i)}\exp((\bm{z}_{i}\cdot\bm{z}_{a})/\tau)}.$ (26)
Theoretical Generalization of Supervised Contrastive Loss. Our formulation (Equation 13) directly generalizes the standard supervised contrastive loss (SupCon) (Khosla et al., 2020), which uniformly weights all positives. When $\mathcal{K}_{i,p}^{s}=1$ and $\mathcal{K}_{i,p}^{d}=1$ (see Equation 17), it reduces exactly to Equation 3. This structural compatibility would be lost if we instead redefined or modulated the temperature.
Role and Effectiveness of Temperature in Contrastive Loss. In the original contrastive learning framework, the temperature hyperparameter $\tau$ controls the sharpness of the softmax distribution over similarity scores. As demonstrated in prior works such as SimCLR (Chen et al., 2020) and further analyzed by Wang & Liu (2021), $\tau$ regulates the concentration level of learned representations: a smaller $\tau$ increases sensitivity to similarity differences, while a larger $\tau$ smooths the gradient signal.
Crucially, temperature $\tau$ acts globally, affecting all pairwise comparisons, including both positive and negative samples. This global effect makes $\tau$ unsuitable for capturing fine-grained, relation-specific weightings such as the similarity–dissimilarity factor, which we intend to apply only to positive samples based on their multi-label relations to the anchor.
From a mathematical perspective, applying the similarity–dissimilarity factor to scaling $\tau$ introduces several issues:
  • Global impact on loss structure: Scaling $\tau$ alters contributions of both positive and negative samples, contradicting our objective of selectively modulating the weights of positive pairs based on semantic relations (R2R2R5R5).
  • Instability of the loss function: Contrastive loss is highly sensitive to variations in $\tau$. Allowing $\tau$ to vary with each pair according to the similarity–dissimilarity factor may lead to unstable gradients, particularly when the factor approaches 0.
  • Theoretical inconsistency: Since $\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}\in[0,1]$ but $\tau\neq 0$ by definition, scaling $\tau$ by this factor is mathematically problematic.
A.2 Choice of dissimilarity penalty function
Motivation for Using $\mathcal{K}^{d}(x)=\frac{1}{1+x}$.
The dissimilarity factor is intended to softly penalize positive samples that contain additional labels not present in the anchor. Our goals in selecting the function $\mathcal{K}^{d}(x)=\frac{1}{1+x}$ were to ensure the following properties:
  • Boundedness: The function maps dissimilarity to the closed interval $(0,1]$, which ensures that the contrastive weight remains numerically stable during training.
  • Monotonic decay: As dissimilarity increases ($x\uparrow$), the penalty decreases smoothly, discouraging high-variance positives from dominating the loss.
  • No hyperparameter sensitivity: The function is parameter-free, making it robust across datasets without the need to tune decay rates or thresholds.
These properties collectively ensure stable gradients, interpretability, and alignment with the goal of penalizing semantically distant positives without completely discarding them.
Comparison to exponential decay function $\mathcal{K}^{d}=\exp(-\alpha x)$:
  • While this offers a sharper decay, it introduces a sensitive hyperparameter $\alpha$
  • For small values of $x$, the decay may be too aggressive or too mild depending on $\alpha$, making behavior inconsistent across datasets with different label cardinalities.
  • It can result in vanishing gradients for large $x$, reducing the contribution of moderately dissimilar positives more than desired.
Simplicity, Interpretability, and Compatibility. The chosen form $\mathcal{K}^{d}(x)=\tfrac{1}{1+x}$ offers several advantages:
  • Simplicity: no tuning required;
  • Smoothness: continuous, differentiable, non-zero everywhere. Because $x\geq 0$ and $\frac{1}{1+x}$ is only non-differentiable when x=1x=-1.
  • Compatibility with our similarity term $K^{s}$, which also scales linearly in set cardinality.
Together, the similarity-dissimilarity factor $\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}$ maintains closed-form interpretability, avoids numerical instability, and allows for clean theoretical analysis, as seen in Theorems 1–5.
Appendix B Theorem Proofs
B.1 Proof of Theorem 1
@@
$|\tilde{\bm{y}}_{p}^{s}|$ $=|\mathcal{S}\cap\mathcal{T}|=|\varnothing|=0$
$\therefore\mathcal{K}_{1}^{s}$ $=\frac{|\tilde{\bm{y}}_{p}^{s}|}{|\tilde{\bm{y}}_{i}|}=\frac{0}{|\tilde{\bm{y}}_{i}|}=0$
@@ B.3 Proof of Theorem 3
We first establish the strict positivity. Given $|\mathcal{S}|>0$ and conditions 29-30, we have:
@@
Case 1 ($m=3$): By Equation 10, we have three conditions: $\mathcal{S}\cap\mathcal{T}\neq\varnothing$, $\mathcal{S}\nsubseteq\mathcal{T}$, and $\mathcal{T}\nsubseteq\mathcal{S}$. These conditions lead to:
@@
Case 2 ($m=4$): When $m=4$, by Equation 11, we have $\mathcal{S}\supseteq\mathcal{T}$. This subset relation implies:
@@
Case 3 ($m=5$): When $m=5$, by Equation 12, we have $\mathcal{S}\subsetneq\mathcal{T}$. This subset relation implies: